Four coins are tossed. Let Y be the random variable representing the number of tails that occur. Find the values of the random variable Y . Construct a probability distribution.

1
Expert's answer
2022-03-01T10:10:12-0500

There are "2^4=16" possible outcomes


"S=\{HHHH, HHHT, HHTH, HTHH, THHH,"

"HHTT, HTHT, HTTH, THTH, THHT, TTHH,"

"HTTT, THTT, TTHT, TTTH, TTTT\}"

The possible values of the random variable "Y" are "0, 1, 2, 3, 4."

We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2"


Construct the probability distribution of the random variable


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 y & 0 & 1 & 2 & 3 & 4 \\ \hline
 p(y) & 1/16 & 1/4 & 3/8 & 1/4 & 1/16
\end{array}"

or


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 y & 0 & 1 & 2 & 3 & 4 \\ \hline
 p(y) & 0.0625 & 0.25 & 0.375 & 0.25 & 0.0625
\end{array}"